Tiki has a total of thirty seven nickels and dimes the dimescome to 40cents more than nickels how many of each does she have
Answer by scott8148(6628)

n + d = 37 ___ n = 37 - d

5n + 40 = 10d

substituting ___ 5(37 - d) + 40 = 10d

185 - 5d + 40 = 10d

225 = 15d ___ 15 = d